OpenStudy (anonymous):

Write an expression in simplest form for the area of a rectangle with a length of 4x and a width of 7y units.

OpenStudy (anonymous):

A=length times width

Directrix (directrix):

@ladya From above, A = l * w A = (4x) * (7y) = ?